OpenStudy (flcon04):

5(6x+4)+1 =-39 Show steps please.

OpenStudy (flcon04):

Another question, Let x be any real number. Then the statement x^3 >0 is true for A. x>0 only B x<0 only C no values of x D all real values of x

OpenStudy (abdullahm):

Welcome to OpenStudy! For the first question, use the distributive property to simplify it first: a(b+c) = ab + ac So, what will 5(6x+4) = ??

OpenStudy (prettygirl_shynice):

il help

OpenStudy (prettygirl_shynice):

Step 1: Simplify both sides of the equation. 5(6x+4)+1=−39 Simplify: (Hide steps) (5)(6x)+(5)(4)+1=−39(Distribute) 30x+20+1=−39 (30x)+(20+1)=−39(Combine Like Terms) 30x+21=−39 30x+21=−39

OpenStudy (prettygirl_shynice):

Step 2: Subtract 21 from both sides. 30x+21−21=−39−21 30x=−60

OpenStudy (prettygirl_shynice):

Step 3: Divide both sides by 30. 30x/30=-60/30 x=?

OpenStudy (canyoupleasehelp):